OpenStudy (anonymous):

x^3 - 7x^2 - 14x + 120 ? Help please :(

OpenStudy (anonymous):

u have to factoring?

OpenStudy (shubhamsrg):

You have to take the derivative of it? (I hope so)

OpenStudy (kainui):

We can't help unless you ask a question I'm afraid.

OpenStudy (anonymous):

Factorise

OpenStudy (anonymous):

synthetic division?

OpenStudy (anonymous):

do you know how to do it?

OpenStudy (anonymous):

I just need to know how to factorise it

OpenStudy (anonymous):

and i am telling you a way of doing it so you dont know synthetic division hmm lets see

OpenStudy (anonymous):

I know how to do it when it's x^2 +x -a just not this one

OpenStudy (anonymous):

yes i got it do you know rational root test?

OpenStudy (anonymous):

she leeeeft sigh

OpenStudy (anonymous):

Let \(\Large f(x)=x^3 - 7x^2 - 14x + 120=0\) \(\Large f(x)\) has 3 roots: \(\Large x=-4,5 \) and \(\Large 6\) Thus, \(\Large x^3 - 7x^2 - 14x + 120=(x+4)(x-5)(x-6)\)

OpenStudy (anonymous):

^ ?

OpenStudy (anonymous):

@aajugdar What?

OpenStudy (anonymous):

That's right, but how did you work out that it's -4, 5 and 6?

OpenStudy (anonymous):

Just use \(\Large Cardano\ Formula\): https://proofwiki.org/wiki/Cardano's_Formula

OpenStudy (anonymous):

idk this formula but lol looks complicated to me

OpenStudy (anonymous):

@BSwan do you know this method?

OpenStudy (anonymous):

yeah , cardano lol just apply the formula but there is much easier way from cardano :O

OpenStudy (anonymous):

factorize 120 to 6,-6,2,-2,4,-4,3,-3,5,-5,24,-24,12,-12,10,-10.. hehe and try f(x) that =0 if ur lucky u would find one of them :) ... x^3 - 7x^2 - 14x + 120=(x-6)(x^2-x-20)